find the measure of an angle whose measure is 26 less than three times the measure of its complement. Thanks & regards. Azhar
Answer by mananth(12270)

Let angle be x
complement will be 990-x)
3 times (90-x)= 3(90-x)
less 26
x=3(90-x)-26
x=270-3x-26
4x=244
x= 61 deg
ananth